# A structure theorem for rooted connectivity in bidirected graphs

### v1

@@ Abstract.
In this paper, we show a structure theorem about rooted connectivity in bidirected graphs in terms of directed graphs. As applications, we prove Lovász’ flame theorem, Pym’s theorem and a strong variant of Menger’s theorem for a class of bidirected graphs and provide counterexamples in the general case.
@@ 1. Introduction
A bidirected graph is a graph $B$ endowed with a sign function $\sigma$ that assigns a sign $\{+,-\}$ to each end of every edge of $B$. The purpose of the sign function is to restrict the structure of walks in the bidirected graph: if two edges uvuv and vwvw appear consecutively in a walk, then the sign of edge uvuv at $v$ must be the opposite of the sign of edge vwvw at $v$. Stated informally: every walk must change signs when passing through a vertex.
Bidirected graphs were introduced independently by Kotzig [Kotzig1959, Kotzig1959b, Kotzig1960] and by Edmonds and Johnson [edmonds2003matching]. Recently, much research on bidirected graphs has been motivated by the fact that they are related to the structure of perfect matchings in (undirected) graphs. In particular, there is a way to translate between undirected graphs with perfect matchings and bidirected graphs. Suppose that $G$ is an undirected graph with a perfect matching $M$. We construct a bidirected graph $(B,\sigma)$ from $G$ as follows. For each edge uvuv of $M$:
@@
  • assign every edge uwuw with $w\neq v$ the sign - at $u$;
  • assign every other edge vwvw with $w\neq u$ the sign ++ at $v$; and
  • contract edge uvuv.
@@
There are several open problems and research directions in structural graph theory dedicated to studying the structure of perfect matchings in graphs. Matching minors and perfect matching width were introduced as a substructure and a width parameter that interact well with the structure of matchings in graphs. Norin proposed the Matching Grid Conjecture, an analog of the Grid Minor Theorem, to relate these two concepts. The conjecture states that, given a graph with a perfect matching, either the graph has bounded perfect matching width or it contains a large grid as a matching minor. Using the correspondence between graphs with perfect matchings and bidirected graphs, one can show that matching minors are a bidirected graph analog of butterfly minors (which are themselves a directed graph analog of minors). In fact, by translating matching minors, perfect matching width, and the Matching Grid Conjecture to the bidirected graph setting, the Matching Grid Conjecture can be understood as a Bidirected Grid Conjecture, i.e. a conjectured grid theorem for bidirected graphs.
Recently, Hatzel, Rabinovich, and Wiederrecht [hatzel2019cyclewidth] proved the Matching Grid Conjecture for bipartite graphs. Their proof relied on two important insights. First, if we start with a bipartite graph $G$ with a perfect matching and construct a bidirected graph $(B,\sigma)$ from $G$ using the procedure described above, the bidirected graph $(B,\sigma)$ is in fact a directed graph, i.e. every edge in $(B,\sigma)$ can be traversed in at most one direction. They then used the Directed Grid Theorem to prove the Matching Grid Conjecture in this case.

### v2

@@ Abstract.
In this paper, we show a structure theorem about rooted connectivity in bidirected graphs in terms of directed graphs. As applications, we prove Lovász’s flame theorem, Pym’s theorem and a strong variant of Menger’s theorem for a class of bidirected graphs and provide counterexamples in the general case.
@@ 1. Introduction
A bidirected graph is a graph $B$ endowed with a sign function $\sigma$ that assigns a sign $\{+,-\}$ to each end of every edge of $B$. The purpose of the sign function is to restrict the structure of walks in the bidirected graph: if two edges $uv$ and $vw$ appear consecutively in a walk, then the sign of edge $uv$ at $v$ must be the opposite of the sign of edge $vw$ at $v$. Stated informally: every walk must change signs when passing through a vertex.
Bidirected graphs were introduced independently by Kotzig [Kotzig1959, Kotzig1959b, Kotzig1960] and by Edmonds and Johnson [edmonds2003matching]. Recently, much research on bidirected graphs has been motivated by the fact that they are related to the structure of perfect matchings in (undirected) graphs [wiederrecht2020thesis, bowler2023mengertheorembidirected, bowler2024hitting, bowler2025ConnectoidsI, KITA2017, Nickel2025, Ghorbani2025]. In particular, there is a way to translate between undirected graphs with perfect matchings and bidirected graphs. Suppose that $G$ is an undirected graph with a perfect matching $M$. We construct a bidirected graph $(B,\sigma)$ from $G$ as follows. For each edge $uv$ of $M$:
@@
  • assign every edge $uw$ with $w\neq v$ the sign - at $u$;
  • assign every other edge $vw$ with $w\neq u$ the sign ++ at $v$; and
  • contract edge $uv$.
@@
There are several open problems and research directions in structural graph theory dedicated to studying the structure of perfect matchings in graphs. Matching minors and perfect matching width were introduced as a substructure and a width parameter that interact well with the structure of matchings in graphs. Norin [norine2005matching] proposed the Matching Grid Conjecture, an analog of the Grid Minor Theorem, to relate these two concepts. The conjecture states that, given a graph with a perfect matching, either the graph has bounded perfect matching width or it contains a large grid as a matching minor. Using the correspondence between graphs with perfect matchings and bidirected graphs, one can show that matching minors are a bidirected graph analog of butterfly minors (which are themselves a directed graph analog of minors). In fact, by translating matching minors, perfect matching width, and the Matching Grid Conjecture to the bidirected graph setting, the Matching Grid Conjecture can be understood as a Bidirected Grid Conjecture, i.e. a conjectured grid theorem for bidirected graphs.
Recently, Hatzel, Rabinovich, and Wiederrecht [hatzel2019cyclewidth] proved the Matching Grid Conjecture for bipartite graphs. Their proof relied on two important insights. First, if we start with a bipartite graph $G$ with a perfect matching and construct a bidirected graph $(B,\sigma)$ from $G$ using the procedure described above, the bidirected graph $(B,\sigma)$ is in fact a directed graph, i.e. every edge in $(B,\sigma)$ can be traversed in at most one direction. They then used the Directed Grid Theorem [kawarabayashi2015directed] to prove the Matching Grid Conjecture in this case.